OpenStudy (clamin):

What are the zeros of the function y= 2x^2+9x+4? a. x= -1/2, x= -4 b. x= -1/2 , x= 4 c x=1/2 , x=4 d. x=1/2, x= -4

OpenStudy (janu16):

A

OpenStudy (janu16):

-1/2 and -4

OpenStudy (marcelie):

it this pre cal?

Directrix (directrix):

Factor this: 2x^2+9x+4 =0 (x+4) * (2x + 1) = 0 x + 4 = 0 or 2x + 1 = 0 Solve each equation for x to get the two roots. @clamin

jimthompson5910 (jim_thompson5910):

Alternative Method: For each answer choice, plug in the given x values one at a time. You can represent `1/2` as `0.5` if fractions throw you off. IF you plug in a certain x value and you get 0 as a result, then that x value you plugged in is a root. a `root` is another term for `zero` or `x-intercept`